OpenStudy (maheshmeghwal9):

Prove that limit of (1+a^x)^1/x @ x tending to 0 is e^a

OpenStudy (maheshmeghwal9):

\[\lim_{x \rightarrow 0}(1+a^x)^(1/x)=e^a\]

OpenStudy (kinggeorge):

\[\lim_{x \rightarrow 0}(1+a^x)^{1/x}=e^a\]

OpenStudy (maheshmeghwal9):

ya !

OpenStudy (zarkon):

not true

OpenStudy (anonymous):

you can try using l'hopital and see

OpenStudy (maheshmeghwal9):

It is 100% true

OpenStudy (maheshmeghwal9):

How?

OpenStudy (zarkon):

really...can't wait to see the proof :)

OpenStudy (anonymous):

i think it is zero

OpenStudy (maheshmeghwal9):

no

OpenStudy (anonymous):

oops wolfram beat me again. depends on the direction

OpenStudy (experimentx):

1/x -> inf a^x = 1

OpenStudy (maheshmeghwal9):

what have u done

OpenStudy (anonymous):

i think maybe the quesion is \[\lim_{x\to 0}(1+ax)^{\frac{1}{x}}\] then you get \(e^a\)

OpenStudy (maheshmeghwal9):

no

OpenStudy (maheshmeghwal9):

but if u prove that one i will do this

OpenStudy (maheshmeghwal9):

ok sorry i think i have wrong question

OpenStudy (anonymous):

imagine

OpenStudy (kinggeorge):

You might want to try finding this limit\[\large \lim_{x \rightarrow \infty} \left(1+{1 \over x}\right)^{ax}\]

OpenStudy (maheshmeghwal9):

ok!will u?

OpenStudy (anonymous):

one line \[e=\large \lim_{x \rightarrow \infty} \left(1+{1 \over x}\right)^{x}\implies e^a=\large \lim_{x \rightarrow \infty} \left(1+{1 \over x}\right)^{ax}\]

OpenStudy (kinggeorge):

I'll get you started at least. First, let\[\large y= \lim_{x \rightarrow \infty} \left(1+{1 \over x}\right)^{ax}\]Now you need to take the natural log of both sides. \[\large \ln(y)= \lim_{x \rightarrow \infty} \ln\left(\left(1+{1 \over x}\right)^{ax}\right)\]Simplify a bit, and then use L'hopitals Rule.

OpenStudy (maheshmeghwal9):

wait please

OpenStudy (experimentx):

Or this should also work out \[ \huge e^{ \lim_{x->0}\ln(1+\frac{1}{x})^{ax}}\]

OpenStudy (experimentx):

then let 1/x = h, h-> inf

OpenStudy (anonymous):

depends on your definition of \(e\) maybe but one definition is \(e=\lim_{x\to \infty}(1+\frac{1}{x})^x\) in which case it is immediate

OpenStudy (maheshmeghwal9):

I can't do that king george please tell me how to do it coz i am only beginning the limit topic

OpenStudy (kinggeorge):

\[\large \ln(y)= \lim_{x \rightarrow \infty} \ln\left(\left(1+{1 \over x}\right)^{ax}\right)\]\[\large \ln(y)= \lim_{x \rightarrow \infty} (ax)\cdot\ln\left(1+{1 \over x}\right)\]This is an indeterminate of the form \(\infty\cdot 0\) so we can use L'hopitals rule on both sides.

OpenStudy (maheshmeghwal9):

k! I try

OpenStudy (kinggeorge):

Good luck.

OpenStudy (maheshmeghwal9):

thanx!

OpenStudy (anonymous):

@kingGeorge \[ \large \ln(y)= \ (ax)\cdot\ln\left(1+{1 \over x}\right)=\\ \frac {\ln\left(1+{1 \over x}\right)} {\frac 1{ax}} \] Apply L'Hospital's rule again.

OpenStudy (maheshmeghwal9):

why did u divide by 1/ax.

OpenStudy (anonymous):

to get the form 0/0 when x tends to infinity

OpenStudy (maheshmeghwal9):

have u taken derivative?

OpenStudy (maheshmeghwal9):

i m getting confused .i don't know anything please solve fully.

OpenStudy (anonymous):

Derivative of Numerator is \[ -\frac{1}{\left(\frac{1}{x}+1\right) x^2} \] Drivative of Denominator is \[ -\frac{1}{a x^2} \] It should be easy from now on.

OpenStudy (anonymous):

The ration of the derivatives after simplifcation becomes \[ \frac{a}{\frac{1}{x}+1} \]

OpenStudy (anonymous):

ratio

OpenStudy (maheshmeghwal9):

k!

OpenStudy (maheshmeghwal9):

I got thanx!